# Training-Free Activation Sparsity in Large Language Models

### v2

@@ 1 Introduction
Large language models (LLMs) demonstrate that scaling in both parameter count and training data leads to capabilities that are useful for addressing a variety of downstream tasks (Brown et al., 2020). However, the large number of parameters in modern LLMs can lead to substantial challenges during inference. In typical small-batch deployment settings, autoregressive inference is memory-bound, i.e., bottlenecked by the speed at which the parameters can be moved from off-chip to on-chip memory. This is in contrast to LLM training and prefill inference, which is generally compute-bound, i.e., bottlenecked by the speed at which computation can performed. A core strategy for overcoming this memory wall (Gholami et al., 2024) is through weight quantization (Frantar et al., 2022; Shao et al., 2023; Yuan et al., 2023; Lin et al., 2024; Dettmers et al., 2023c; Tseng et al., 2024; Egiazarian et al., 2024) and sparsification (Wang et al., 2019; Frantar & Alistarh, 2023; Xia et al., 2023; Ma et al., 2023), which can lead to practical speed-ups when coupled with specialized kernels that move the weights from off-chip to on-chip memory in quantized/sparse formats (Kim et al., 2023; Dettmers et al., 2023b; Frantar et al., 2024; Wang et al., 2024b; Xia et al., 2024; Guo et al., 2024).
The above methods directly compress a model’s weights and apply the same (quantized/sparse) matrix to all inputs. Activation sparsity (Chen et al., 2023; Raihan & Aamodt, 2020; Kurtz et al., 2020) is an alternative method which enforces input-dependent structure on the weight matrices by leveraging (or inducing) sparsity in the hidden states. Since the weight channels corresponding to zero-valued activations are not used in computation, speed-up can be realized by selectively omitting these weights during memory transfer, which is possible due to the hardware-friendly channel-wise sparsity pattern. In older LLMs, activation sparsity is largely made possible by the high natural sparsity (around 95%) in the intermediate states of the MLP blocks in ReLU-based Transformer models (Li et al., 2023). Based on this, Liu et al. (2023) propose DejaVu, which learns a small auxiliary model that predicts the contextual activation sparsity patterns of future layers, and realize a 2×2$\times$2 × wall-clock speed-up on OPT-175B (Zhang et al., 2022). Because the hidden state is extremely sparse, the less expressive auxiliary model can afford to overestimate non-zero activations while maintaining accuracy and efficiency (e.g., 20% predicted vs. 5% actual non-zero entries).
@@ 2 Related Work
Conditional computation (Bengio, 2013; Bengio et al., 2016) alleviates the burden of training and serving by selectively activating parts of a model. Shazeer et al. (2017) propose Mixture-of-Experts (MoE) in language models, applying conditional computation to the feed forward networks. Mixture-of-Experts models decouple parameter count with computational footprint (Fedus et al., 2022), and demonstrate superior scaling laws compared to dense baselines (Clark et al., 2022).
@@
Lee et al. (2024a) propose CATS, and realize training-free activation sparsity on SwiGLU based LLMs by applying magnitude pruning on the output of $\textbf{W}_{\text{gate}}$, with the intuition that in the training-free setting, ReLU-based methods suboptimally zero out nontrivial negative values but keep positive values with lower magnitude intact. They achieve up to 50% sparsity in $\textbf{W}_{\text{up}}$ and $\textbf{W}_{\text{down}}$ for Mistral and Llama-2-7B. However, other matrices including $\textbf{W}_{\text{gate}}$ and $\textbf{W}_{\text{q,k,v,o}}$ are computed densely, resulting in lower model-wide sparsity (roughly 25%), whereas we target every matrix in the model.
@@ 3 Background: Activation Sparsity in Neural Networks
The activation sparsity of a hidden state $\mathbf{x}$ is defined as the proportion of zero-valued entries, which can interact with the model in two ways. The first is input sparsity: when computing $\mathbf{y}=\mathbf{x}\mathbf{W}^{\top}$ for $\mathbf{x}\in\mathbb{R}^{m},\mathbf{W}\in\mathbb{R}^{n\times m}$, the columns $\mathbf{W}_{:,i}$ corresponding to zero-valued entries $\mathbf{x}_{i}$ are unused. The second is output sparsity: when computing $\mathbf{y}=\mathbf{s}\odot(\mathbf{x}\mathbf{W}^{\top})$ for the aforementioned parameters and mask $\mathbf{s}\in\mathbb{R}^{n}$, the rows $\mathbf{W}_{i,:}$ corresponding to zero-valued entries $\mathbf{s}_{i}$ are unused. Recent work makes use of output sparsity on GLU variants, treating $\mathbf{s}=\mathrm{sparsify}(\sigma(\mathbf{x}\mathbf{W}_{\text{gate}}^{\top}))$ as the mask and applying output sparsity on $\mathbf{xW}_{\text{up}}^{\top}$, with the intuition that $\sigma(\cdot)$ serves as a gating mechanism. Interestingly, we find in Section  5.4.1 that input sparsity is still preferable in the training-free case for SwiGLU.
@@ 4.1 Motivating Study: Distributional Properties of Activations in LLMs
We perform a preliminary study of the distributional properties of activations of LLMs. We collect activations of Llama-3-8B (Dubey et al., 2024) sampled from C4 (Dodge et al., 2021) at the four hidden states in a Transformer block, [^1] Throughout, we use “block” to refer to an entire Transformer block consisting of the seven matrices and “layer” to refer to an individual layer (corresponding to a single matrix) within the Transformer block. and visualize them in Figure 2. As indicated by prior work, some of the activations are heavy-tailed and contain outliers (Dettmers et al., 2022; Xiao et al., 2022; Wei et al., 2022; Nrusimha et al., 2024). The hidden states are moreover zero-mean unimodal, and qualitatively fall into two distinctly shaped distributions. The hidden states before the Attention and the MLP layers tend to be Gaussian-like, while the hidden states in the intermediate of such layers tend to be Laplacian-like. The concentration of the activations around zero motivates our magnitude-based activation pruning approach.
@@ 4.2 TEAL
# Find size (memory footprint) of matrices
@@
# Init block and all layer sparsities to zero
@@
     $p_{j}\mathrel{+}=\delta_{j}$
@@ 4.3 Block-wise Greedy Optimization
For each Transformer block, we seek to minimize the block-wise $\ell_{2}$ activation error subject to a block-level sparsity constraint. Each Transformer block consists of seven matrices: $\mathbf{W}_{\text{q}},\mathbf{W}_{\text{k}},\mathbf{W}_{\text{v}},\mathbf{W}_{\text{o}},\mathbf{W}_{\text{gate}},\mathbf{W}_{\text{up}},\mathbf{W}_{\text{down}}$. Algorithm 1 initializes the sparsity levels of all layers to zero, and attempts to increment the sparsity level of each layer by an amount inversely proportional to its memory footprint. The layer with the lowest $\ell_{2}$ activation error is incremented, and the block-level sparsity plus associated layer-level sparsities are recorded. We assign the same block-level sparsity level across all Transformer blocks; therefore, all blocks have the same target sparsity level, but the individual layer-level sparsities could be different across different blocks.
@@ 4.4 Hardware Aware Acceleration
Our kernel makes the following improvements on top of the original kernel: (1) We fuse the mask creation process, as $\mathbf{s}=\mathbf{x}[|\mathbf{x}|>t_{p}]$ is entirely characterized by $\mathbf{x}$ and $t_{p}$ in TEAL; (2) We accumulate along the outer SplitK dimension in FP16 (keeping the inner in-register accumulation in FP32), as writing to global memory in FP32 results in significant traffic; (3) We specify an eviction policy in NVIDIA PTX, prioritizing cache retention for activations which are reused across multiple thread blocks, and deprioritizing weights which are block-specific. This guarantees that activations are persistent in L2 cache.
Figure 3: Latency vs. sparsity for matrix-vector multiplication (1x4096 × 4096x14336), comparing TEAL to Deja Vu.
@@ 5 Results
We use the block-wise greedily optimized sparsities from Section 4.3 for TEAL, and primarily compare to CATS (Lee et al., 2024a) and ReLUfication (Mirzadeh et al., 2023) in their training-free configurations with no finetuning. We report model-level sparsities for all methods.
Both CATS and ReLUfication apply sparsity to MLP parameters, and do not apply sparsity to attention parameters. For SwiGLU variants, the original dense forward pass is defined as: $\text{MLP}(\mathbf{x})=(\text{SiLU}(\mathbf{x}\mathbf{W}_{\text{gate}})\odot\mathbf{x}\mathbf{W}_{\text{up}})\mathbf{W}_{\text{down}}$. CATS sparsifies the output of $\mathbf{W}_{\text{gate}}$, replacing $\text{SiLU}(\mathbf{x}\mathbf{W}_{\text{gate}})$ with $s_{t_{p}}(\text{SiLU}(\mathbf{x}\mathbf{W}_{\text{gate}}))$ for sparsification function $s_{t_{p}}$ associated with the distribution of $\text{SiLU}(\mathbf{x}\mathbf{W}_{\text{gate}})$. ReLUfication simply replaces the SiLU activation function with ReLU. Both methods end up sparsifying the intermediate state of the MLP by first performing dense computation on $\mathbf{W}_{\text{gate}}$, enforcing output sparsity on $\mathbf{W}_{\text{up}}$, and then enforcing input sparsity on $\mathbf{W}_{\text{down}}$. This is in contrast with TEAL, which enforces input sparsity on all matrices.
@@ 5.1 Accuracy
Table 2: Downstream task evaluation results. Reported results are averaged over six tasks. See Appendix A.2 for fine-grained results. We omit CATS 40% and ReLUfication as they are degenerate.
@@ 5.2 End-to-end Decoding Speed-up
We benchmark TEAL’s end-to-end single-batch decoding latency by integrating it with GPT-Fast (PyTorch, 2024). We enable CUDA graphs and torch.compile. Tests use Llama-2 (7B, 13B) and Llama-3 (8B, 70B) models at 0%, 25%, 40%, and 50% uniform sparsities. Our GPU power limit settings are 500W and 300W for A100 and A6000 respectively. As shown in Table 3, TEAL achieves significant speed-ups of up to 1.53×1.53$\times$1.53 × and 1.8×1.8$\times$1.8 × at 40% and 50% sparsity respectively. TEAL is slower than the baseline at 0% sparsity on A100 due to torch.compile strengthening the torch.matmul baseline. This suggests further room for optimization of our kernel. We find lower speedups for Llama-3-8B compared to Llama-2-7B partially due to its larger LM Head, which we do not currently sparsify.

### v3

@@ 1 Introduction
Large language models (LLMs) demonstrate that scaling in both parameter count and training data leads to capabilities that are useful for addressing a variety of downstream tasks (Brown et al., 2020). However, the large number of parameters in modern LLMs can lead to substantial challenges during inference. In typical small-batch deployment settings, autoregressive inference is memory-bound, i.e., bottlenecked by the speed at which the parameters can be moved from off-chip to on-chip memory. This is in contrast to LLM training and prefill inference, which is generally compute-bound, i.e., bottlenecked by the speed at which computation can performed. A core strategy for overcoming this memory wall (Gholami et al., 2024) is through weight quantization (Frantar et al., 2022; Shao et al., 2023; Yuan et al., 2023; Lin et al., 2024; Dettmers et al., 2023c; Tseng et al., 2024; Egiazarian et al., 2024; Liu et al., 2024) and sparsification (Wang et al., 2019; Frantar & Alistarh, 2023; Xia et al., 2023; Ma et al., 2023), which can lead to practical speed-ups when coupled with specialized kernels that move the weights from off-chip to on-chip memory in quantized/sparse formats (Kim et al., 2023; Dettmers et al., 2023b; Frantar et al., 2024; Wang et al., 2024b; Xia et al., 2024; Guo et al., 2024).
The above methods directly compress a model’s weights and apply the same (quantized/sparse) matrix to all inputs. Activation sparsity (Chen et al., 2023; Raihan & Aamodt, 2020; Kurtz et al., 2020) is an alternative method which enforces input-dependent structure on the weight matrices by leveraging (or inducing) sparsity in the hidden states. Since the weight channels corresponding to zero-valued activations are not used in computation, speed-up can be realized by selectively omitting these weights during memory transfer, which is possible due to the hardware-friendly channel-wise sparsity pattern. In older LLMs, activation sparsity is largely made possible by the high natural sparsity (around 95%) in the intermediate states of the MLP blocks in ReLU-based Transformer models (Li et al., 2023). Based on this, Liu et al. (2023) propose DejaVu, which learns a small auxiliary model that predicts the contextual activation sparsity patterns of future layers, and realize a 2×2$\times$2 × wall-clock speed-up on OPT-175B (Zhang et al., 2022a). Because the hidden state is extremely sparse, the less expressive auxiliary model can afford to overestimate non-zero activations while maintaining accuracy and efficiency (e.g., 20% predicted vs. 5% actual non-zero entries).
@@ 2 Related Work
Conditional computation (Bengio, 2013; Bengio et al., 2016) alleviates the burden of training and serving by selectively activating parts of a model. Shazeer et al. (2017) propose Mixture-of-Experts (MoE) in language models, applying conditional computation to feed forward networks. Mixture-of-Experts models decouple parameter count with computational footprint (Fedus et al., 2022), and demonstrate superior scaling laws compared to dense baselines (Clark et al., 2022). Dense models can also be converted into MoE models after pre-training (Zhang et al., 2022b; Szatkowski et al., 2024; Zheng et al., 2024).
@@
Lee et al. (2024a) propose CATS, and realize training-free activation sparsity on SwiGLU based LLMs by applying magnitude pruning on the output of $\textbf{W}_{\text{gate}}$, with the intuition that in the training-free setting, ReLU-based methods suboptimally zero out nontrivial negative values but keep positive values with lower magnitude intact. They achieve up to 50% sparsity in $\textbf{W}_{\text{up}}$ and $\textbf{W}_{\text{down}}$ for Mistral and Llama-2-7B. However, other matrices including $\textbf{W}_{\text{gate}}$ and $\textbf{W}_{\text{q,k,v,o}}$ are computed densely, resulting in lower model-wide sparsity (roughly 25%), whereas we target every matrix in the model. We refer the reader to Appendix A.3 for formal definitions of the weight matrices and their interactions within each Transformer block.
@@ 3 Background: Activation Sparsity in Neural Networks
The activation sparsity of a hidden state $\mathbf{x}$ is defined as the proportion of zero-valued entries, which can interact with the model in two ways. The first is input sparsity: when computing $\mathbf{y}=\mathbf{x}\mathbf{W}^{\top}$ for $\mathbf{x}\in\mathbb{R}^{m},\mathbf{W}\in\mathbb{R}^{n\times m}$, the columns $\mathbf{W}_{:,i}$ corresponding to zero-valued entries $\mathbf{x}_{i}$ are unused. The second is output sparsity: when computing $\mathbf{y}=\mathbf{s}\odot(\mathbf{x}\mathbf{W}^{\top})$ for the aforementioned parameters and mask $\mathbf{s}\in\mathbb{R}^{n}$, the rows $\mathbf{W}_{i,:}$ corresponding to zero-valued entries $\mathbf{s}_{i}$ are unused. CATS makes use of output sparsity on GLU variants, treating $\mathbf{s}=\mathrm{sparsify}(\sigma(\mathbf{x}\mathbf{W}_{\text{gate}}^{\top}))$ as the mask and applying output sparsity on $\mathbf{xW}_{\text{up}}^{\top}$, with the intuition that $\sigma(\cdot)$ serves as a gating mechanism. Interestingly, we find in Section  5.4.1 that input sparsity is still preferable in the training-free case for SwiGLU.
@@ 4.1 Motivating Study: Distributional Properties of Activations in LLMs
We perform a preliminary study of the distributional properties of activations of LLMs. We collect activations of Llama-3-8B (Dubey et al., 2024) sampled from C4 (Raffel et al., 2023) at the four hidden states in a Transformer block, [^1] Throughout, we use “block” to refer to an entire Transformer block consisting of the seven matrices and “layer” to refer to an individual layer (corresponding to a single matrix) within the Transformer block. and visualize them in Figure 2. As indicated by prior work, some of the activations are heavy-tailed and contain outliers (Dettmers et al., 2022; Xiao et al., 2022; Wei et al., 2022; Nrusimha et al., 2024). The hidden states are moreover zero-mean unimodal, and qualitatively fall into two distinctly shaped distributions. The hidden states before the Attention and the MLP layers tend to be Gaussian-like, while the hidden states in the intermediate of such layers tend to be Laplacian-like. The concentration of the activations around zero motivates our magnitude-based activation pruning approach.
@@ 4.2 TEAL
# Record size (memory footprint) of matrices
@@
# Initialize block and layer sparsities to zero
@@
     $p_{j}\mathrel{+}=\delta_{j}$ Increment layer with lowest error
@@ 4.3 Block-wise Greedy Optimization
For each Transformer block, we seek to minimize the block-wise $\ell_{2}$ activation error subject to a block-level sparsity constraint. Each Transformer block consists of seven matrices: $\mathbf{W}_{\text{q}},\mathbf{W}_{\text{k}},\mathbf{W}_{\text{v}},\mathbf{W}_{\text{o}},\mathbf{W}_{\text{gate}},\mathbf{W}_{\text{up}},\mathbf{W}_{\text{down}}$. Algorithm 1 initializes the sparsity levels of all layers to zero, and attempts to increment the sparsity level of each layer by an amount inversely proportional to its memory footprint. The layer with the lowest $\ell_{2}$ activation error is incremented, and the block-level sparsity plus associated layer-level sparsities are recorded. We assign the same block-level sparsity level to every Transformer block; therefore, all blocks have the same target sparsity level, but the individual layer-level sparsities could be different across different blocks.
@@ 4.4 Hardware Aware Acceleration
Our kernel makes the following improvements on top of the original kernel: (1) We fuse the mask creation process, as $\mathbf{s}=\mathbf{x}[|\mathbf{x}|>t_{p}]$ is entirely characterized by $\mathbf{x}$ and $t_{p}$ in TEAL; (2) We accumulate along the outer SplitK dimension in FP16 (keeping the inner in-register accumulation in FP32), as writing to global memory in FP32 results in significant traffic; (3) We specify an eviction policy in PTX, prioritizing cache retention for activations which are reused across multiple thread blocks, and deprioritizing weights which are block-specific. This guarantees that activations are persistent in L2 cache.
Figure 3: Latency vs. sparsity for matrix-vector multiplication (1x4096 × 4096x14336), comparing TEAL to Deja Vu. ’Theoretical Optimal’ shows the latency reduction for torch.matmul assuming perfect linear scaling with sparsity.
@@ 5 Results
We use the block-wise greedily optimized sparsities from Section 4.3 for TEAL, and primarily compare to CATS (Lee et al., 2024a) in its training-free configuration with no finetuning. We report model-level sparsities for all methods.
CATS applies sparsity to MLP parameters, and does not apply sparsity to attention parameters. In particular, CATS sparsifies the output of $\mathbf{W}_{\text{gate}}$, replacing $\text{SiLU}(\mathbf{x}\mathbf{W}_{\text{gate}})$ with $s_{t_{p}}(\text{SiLU}(\mathbf{x}\mathbf{W}_{\text{gate}}))$ for sparsification function $s_{t_{p}}$ associated with the distribution of $\text{SiLU}(\mathbf{x}\mathbf{W}_{\text{gate}})$. Overall, CATS sparsifies the intermediate state of the MLP by first performing dense computation on $\mathbf{W}_{\text{gate}}$, enforcing output sparsity on $\mathbf{W}_{\text{up}}$, and then enforcing input sparsity on $\mathbf{W}_{\text{down}}$. This is in contrast with TEAL, which enforces input sparsity on all matrices.
@@ 5.1 Accuracy
Table 2: Downstream task evaluation results. Reported results are averaged over six tasks. See Appendix A.2 for fine-grained results. We omit CATS 40% as it is degenerate.
@@ 5.2 End-to-end Decoding Speed-up
We benchmark TEAL’s end-to-end single-batch decoding latency by integrating it with GPT-Fast (PyTorch, 2024). We enable CUDA graphs and torch.compile. Tests use Llama-2 (7B, 13B) and Llama-3 (8B, 70B) models at 0%, 25%, 40%, and 50% uniform sparsities. We use the standard inference benchmarking setup in GPT-Fast, which passes in roughly 5 input tokens and generates at most 200 output tokens. Our GPU power limit settings are 500W and 300W for A100 and A6000 respectively. As shown in Table 3, TEAL achieves significant speed-ups of up to 1.53×1.53$\times$1.53 × and 1.8×1.8$\times$1.8 × at 40% and 50% sparsity respectively. TEAL is slower than the baseline at 0% sparsity on A100 due to torch.compile strengthening the torch.matmul baseline. This suggests further room for optimization of our kernel. We find lower speedups for Llama-3-8B compared to Llama-2-7B partially due to its larger LM Head, which we do not currently sparsify. We leave the sparsification of LM Head to future work.
@@ 7 Conclusion
Acknowledgments
We thank Neil Movva, Jue Wang, and Yucheng Lu for helpful comments and discussion.
@@ A.2 Full Downstream Task Results
A.3 Transformer Architecture Overview
A Transformer block consists of an attention layer followed by a multilayer perceptron (MLP). Each block contains seven weight matrices that process the input $\mathbf{x}\in\mathbb{R}^{d}$ in sequence:
Attention:
In Grouped Query Attention (GQA) (Ainslie et al., 2023), the matrices $\mathbf{W}_{\text{q}}\in\mathbb{R}^{d\times d}$ and $\mathbf{W}_{\text{k}},\mathbf{W}_{\text{v}}\in\mathbb{R}^{d_{h}\times d}$ project the input into query, key, and value representations, which are fed into the attention operation. After the attention operation, $\mathbf{W}_{\text{o}}\in\mathbb{R}^{d\times d_{h}}$ projects the output back to the model dimension.
MLP:
The SwiGLU (Shazeer, 2020) MLP variant uses three matrices: $\mathbf{W}_{\text{gate}},\mathbf{W}_{\text{up}}\in\mathbb{R}^{d_{m}\times d}$ which project to a higher dimension, and $\mathbf{W}_{\text{down}}\in\mathbb{R}^{d\times d_{m}}$ which projects back to the model dimension. The computation flow is:
$\text{MLP}(\mathbf{x})=(\text{SiLU}(\mathbf{x}\mathbf{W}_{\text{gate}}^{\top})\odot\mathbf{x}\mathbf{W}_{\text{up}}^{\top})\mathbf{W}_{\text{down}}^{\top}$
where $\text{SiLU}(\mathbf{x})=\mathbf{x}\odot\sigma(\mathbf{x})$, $\sigma$ is the sigmoid function, and $\odot$ denotes element-wise multiplication.
A.4 Comparison to 2:4 Weight Sparsity
We compare TEAL to MaskLLM (Fang et al., 2024), a state-of-the-art approach to semi-structured 2:4 weight sparsity that learns weight masks through differentiable relaxation. The authors have not released checkpoints as of the time of writing, so we trained our own masks on Llama-3-8B using 2B tokens from C4. We use the greedily optimized sparsities described in Section 4.3 for TEAL, and evaluate both methods on WikiText:
Model Perplexity
Llama-3-8B (baseline) 5.870
Llama-3-8B + TEAL (50%) 6.673
Llama-3-8B + MaskLLM (2:4) 8.532
Llama-3-8B + TEAL + MaskLLM 9.590
We observe that TEAL outperforms MaskLLM, while being training-free. In contrast, MaskLLM is computationally expensive—training on Llama-3-8B requires 2B tokens and approximately 8B frozen parameters plus 12B learnable parameters, which took us roughly 800 H100 hours. The combination of both methods works well, suggesting they are complementary rather than mutually exclusive.
We additionally compare single-batch decoding speed-up on Llama-2-7B using a single A6000 GPU (using MaskLLM’s reported numbers from their Table 6):
Model Speed-up
Llama-2-7B + TEAL (50%) 1.78×
Llama-2-7B + MaskLLM (2:4) 1.4×
This comparison may not be fully representative as 2:4 weight sparsity is more performant in high-batch settings and can additionally accelerate the prefill phase.
A.5 Compatibility with Fine-tuning
Table 10: Perplexity results with and without fine-tuning on Llama-3-8B.
Sparsity Level PPL (No Fine-tuning) PPL (With Fine-tuning)
0% (baseline) 5.870
50% 6.673 6.622
60% 7.827 7.515
70% 13.39 9.927
90% 4.141 $\cdot$ 105 4.589 $\cdot$ 103
While TEAL is primarily designed as a training-free method, it can be further enhanced with fine-tuning. We fine-tune Llama-3-8B using LoRA (Hu et al., 2021) with a rank of 32 (approximately 1% of parameters are trainable) and a learning rate of 0.0002. The model is fine-tuned on 30M tokens from C4. We evaluate on WikiText and use the greedily optimized sparsities described in Section 4.3.
We observe in Table 10 that fine-tuning provides marginal improvements at lower sparsity levels (50-60%). The benefits are more pronounced at higher sparsity levels (70-90%), where fine-tuning helps to recover some of the performance lost due to aggressive sparsification.